OpenStudy (anonymous):

Determine the position and nature of the stationary points of y=ƒ(x) on the interval o≤x≤2π and show them on a sketch graph of the function:

OpenStudy (anonymous):

anything else go with this problem?

OpenStudy (anonymous):

yes sorry c) y=sin^2x

OpenStudy (anonymous):

sorry, i still don't understand the problem. is there some kind of transformation? (reflection, rotation, translation, etc?)

OpenStudy (anonymous):

Determine the position and nature of the stationary points of y=ƒ(x) on the interval o≤x≤2π and show them on a sketch graph of the function: y=sin^2x No. Its on Derivatives of Trigonometric Functions

OpenStudy (anonymous):

I need to solve c.